Gallium has two naturally occurring isotopes. One of them, gallium-69, has a mass of 68.925581 u and a percent isotopic abundance of 60.11%. What must be the mass and percent isotopic abundance of the other isotope, gallium-71?
Answer #1

Ga - 69                        68.925581 amu

abundance                      60.11%

Ga - 71                       

abundance                  100-60.11   =39.89%

atomic mass Ga   = 69.723amu

average atomic mass    =   mass of 1st isotope* abundance + mass of 2nd isotope* abundance/100

69.723                         = 68.925581*60.11 + x*39.89/100

69.723*100                = 68.925581*60.11 + x*39.89

x    = 70.9247amu

The atomic mass of gallium -71 = 70.9247amu